# Fidelity-Imposed Displacement Editing for the Learn2Reg 2024 SHG-BF Challenge

### v1

@@ Abstract
Co-examination of second-harmonic generation (SHG) and bright-field (BF) microscopy enables the differentiation of tissue components and collagen fibers, aiding the analysis of human breast and pancreatic cancer tissues. However, large discrepancies between SHG and BF images pose challenges for current learning-based registration models in aligning SHG to BF. In this paper, we propose a novel multi-modal registration framework that employs fidelity-imposed displacement editing to address these challenges. The framework integrates batch-wise contrastive learning, feature-based pre-alignment, and instance-level optimization. Experimental results from the Learn2Reg COMULISglobe SHG-BF Challenge validate the effectiveness of our method, securing the 1st place on the online leaderboard.
@@ 2 Related Work
Descriptor Matching Learning. Descriptor matching methods [16, 17, 18, 19] effectively address the high foreground-background imbalance by relying on feature descriptors and feature matching, utilizing both handcrafted and learning-based descriptors. Traditional RANSAC [20] and modern graph- and attention-based approaches [21] are commonly used for matching. However, these approaches lack a fidelity loss, leaving the output predictions unconstrained.
@@ 3.1 Features prealignment with contrastive learning
Feature-based prealignment. With the learned representations $C_{A}$ and $C_{B}$, we perform feature-based prealignment. We detect keypoints and extract descriptors using a method inspired by XFeat [24], adapted to our context. We match features between $C_{A}$ and $C_{B}$ to compute an initial affine transformation $\Phi_{\text{init}}$. We also demonstrated the matching keypoints in Fig. 3 in blue dots and connected with green lines.
This prealignment step reduces large discrepancies between the images, providing a good starting point for the subsequent optimization.
@@ 3.2 Test-time Instance Optimization
We adopt the core principles of the ConvexAdam method [25], enabling flexible instance optimization between affine and deformable transformations. Unlike the original ConvexAdam, which optimizes flow parameters using hand-crafted features, our approach leverages learned features from the preceding stage and employs a pyramid structure for multi-resolution alignment across scales $1,\frac{1}{2},\frac{1}{4},\frac{1}{8},\frac{1}{16}$. This hierarchical setup allows for efficient alignment at both large and fine scales.
Following the prealignment performed during the training stage, we use a convolutional network with large kernel sizes to generate the displacement fields. Displacement assignment is achieved by evaluating patch similarity costs and identifying the minimal cost positions. Smooth regularization is then applied through Gaussian smoothing, as described in [26]. Our fidelity loss combines Local Normalized Cross-Correlation (LNCC), which handles local similarities to address the foreground-background imbalance, with Cross Mutual Information Function (CMIF), which considers global intensity relationships to bridge the modality gap.
Since the primary features in SHG images are sparsely distributed, we avoid using the deformation field directly. Instead, we sample displacements based on SHG intensity, followed by least-squares analysis on these sparse displacements. This results in an over-determined system of equations that we solve to obtain the final affine transformation. This approach allows for a smooth transition between deformable and affine transformations, enhancing the precision of the registration results.
@@ 4.1 Dataset
We evaluated our method on the dataset from the Learn2Reg 2024 Challenge Task 3: COMULISglobe SHG-BF [1], which includes 156 training, 10 validation, and 40 test cases of SHG and H&E-stained BF microscopy images of breast and pancreatic cancer tissues. Acquired at the University of Wisconsin-Madison, these images present significant challenges due to differences in imaging modalities: SHG images emphasize collagen fibers, while BF images highlight H&E-stained tissue structures. Reliable expert annotations provide landmarks for validation and testing.
@@ 4.2 Data Preprocessing and Evaluation Metrics
We adopted the evaluation metrics established by the Learn2Reg 2024 Challenge, specifically using Target Registration Error (TRE) as the primary metric to assess registration accuracy. TRE measures the average Euclidean distance between corresponding anatomical landmarks in the fixed and moving images after registration. Lower TRE values indicate better alignment. All reported results on the validation set are based on the challenge leaderboard.
@@ 4.3 Quantitative Results
To evaluate the effectiveness of our proposed components, we conducted ablation studies summarized in Table 1 (second part). Introducing instance-level optimization using LNCC reduced the TRE by approximately 60%, effectively mitigating the imbalance caused by sparse SHG features. Using only CMIF as the optimization metric increased the TRE, suggesting that CMIF alone may lead to local minima due to its global nature and the modality gap. Combining CMIF with LNCC in our fidelity loss improved performance, reducing the TRE from 2.361 to 1.943. This demonstrates that our fidelity-imposed displacement editing, which balances global and local alignment, enhances registration accuracy.
Additionally, we tested different iteration counts during instance-level optimization (15, 30, and 50 iterations) using both CMIF and LNCC (Fig. 5). Our method consistently achieved low TRE values, with the best performance at 30 iterations (mean TREs of 2.759, 1.943, and 2.168, respectively). This indicates our approach’s stability without extensive computation.

### v3

@@ Abstract
Co-examination of second-harmonic generation (SHG) and bright-field (BF) microscopy enables the differentiation of tissue components and collagen fibers, aiding the analysis of human breast and pancreatic cancer tissues. However, large discrepancies between SHG and BF images pose challenges for current learning-based registration models in aligning SHG to BF. In this paper, we propose a novel multi-modal registration framework that employs fidelity-imposed displacement editing to address these challenges. The framework integrates batch-wise contrastive learning, feature-based pre-alignment, and instance-level optimization. Experimental results from the Learn2Reg COMULISglobe SHG-BF Challenge validate the effectiveness of our method, securing the 1st place on the online leaderboard. Our code is available at https://github.com/JackyWang22/Neon_ShgBf.
**footnotetext: Corresponding author. hz459@cornell.edu
@@ 2 Related Work
Descriptor Matching Learning. Descriptor matching methods [16, 17, 18, 19] effectively address the high foreground-background imbalance by relying on feature descriptors and feature matching, utilizing both handcrafted and learning-based descriptors. Traditional RANSAC and modern graph- and attention-based approaches [20] are commonly used for matching. However, these approaches lack a fidelity loss, leaving the output predictions unconstrained.
@@ 3.1 Features prealignment with contrastive learning
Feature-based prealignment. With the learned representations $C_{A}$ and $C_{B}$, we perform feature-based prealignment. We detect keypoints and extract descriptors using a method inspired by XFeat [23], adapted to our context. We match features between $C_{A}$ and $C_{B}$ (Fig. 3 second row) to compute an initial affine transformation $\Phi_{\text{init}}$. We also demonstrated the matching keypoints in Fig. 3 in blue dots and connected with green lines.
@@ 3.2 Test-time Instance Optimization
To address misalignment caused by local contrast noise in contrastive feature mapping, we adopt the core principles of the ConvexAdam method [24], enabling flexible instance optimization between affine and deformable transformations. Unlike the original ConvexAdam, which optimizes flow parameters using hand-crafted features, our approach leverages learned features from the preceding stage and employs a multi-resolution pyramid across scales of $1,\frac{1}{2},\frac{1}{4},\frac{1}{8},\frac{1}{16}$. This hierarchical setup allows for efficient alignment at both large and fine scales.
Following the prealignment, we generate displacement fields using a convolutional network. Displacements are assigned by evaluating patch similarity costs and identifying minimal cost positions. Smooth regularization is then applied through Gaussian smoothing, as described in [25]. Our fidelity loss combines Local Normalized Cross-Correlation (LNCC), which handles local similarities and addresses the foreground-background imbalance, with Cross Mutual Information Function (CMIF), which considers global intensity relationships to bridge the modality gap.
Since the primary features in SHG images are sparsely distributed, we avoid using the deformation field directly. Instead, we sample displacements based on SHG intensity and perform least-squares analysis on these sparse displacements, resulting in an over-determined system of equations that we solve to obtain the final affine transformation. This approach allows for a smooth transition between deformable and affine transformations, enhancing the precision of the registration results.
@@ 4.1 Dataset
We evaluated our method on the dataset from the Learn2Reg 2024 Challenge Task 3: COMULISglobe SHG-BF [1], which includes 156 training, 10 validation, and 40 test cases of SHG and H&E-stained BF microscopy images of breast and pancreatic cancer tissues shown in Fig. 4. Acquired at the University of Wisconsin-Madison, these images present significant challenges due to differences in imaging modalities: SHG images emphasize collagen fibers, while BF images highlight H&E-stained tissue structures. Reliable expert annotations provide landmarks for validation and testing.
@@ 4.2 Data Preprocessing and Evaluation Metrics
We used the Learn2Reg 2024 Challenge metrics, primarily Target Registration Error (TRE), to measure registration accuracy. TRE calculates the average distance between matched landmarks in the fixed and moving images—lower values mean better alignment. Validation results are from the challenge leaderboard.
@@ 4.3 Quantitative Results
To evaluate the effectiveness of our proposed components, we conducted ablation studies summarized in Table 1 (second part). LNCC is less sensitive to local noise and effectively handled the sparse SHG features, reducing the TRE by approximately 60%. However, it is less capable of capturing global structure. In contrast, CMIF captures global structure better but is more sensitive to local noise introduced by the contrastive feature mapping, leading to worse results when used alone. By combining LNCC with CMIF in our fidelity loss, we balanced local robustness and global alignment, reducing the TRE from 2.361 to 1.943. This demonstrates that our fidelity-imposed displacement editing enhances registration accuracy by integrating the strengths of both metrics.
Additionally, we tested different iteration counts during(15, 30, and 50 iterations) using both CMIF and LNCC (Fig. 5). Our method consistently achieved low TRE values, with the best performance at 30 iterations (mean TREs of 2.759, 1.943, and 2.168, respectively). This indicates our approach’s stability without extensive computation.
@@ 5 Conclusion
Acknowledgment. This work is supported, in part, by the National Key Research and Development Program of China under Grant 2022YFE0134700, National Natural Science Foundation of China under Grant U22B2050 and 62425305, and the Science and Technology Program of Changsha under Grant kq2102009. Compliance with Ethical Standards. Ethical approval was not required as confirmed by the license attached with the open access data.